# Recursive Gaussian Process State Space Model

### v3

@@ Abstract
Nomenclature
$\mathcal{GP}$  = Gaussian process (GP)
$m_{0}(\cdot)$  = prior mean function of the GP
$K_{0}(\cdot,\cdot;\bm{\theta})$  = prior covariance function (or kernel) of the GP
$\bm{\theta}$  = kernel hyperparameters, nondimensional unit
$\bm{K}_{ab}$  = kernel matrix indexed by function outputs or function inputs, nondimensional unit
$\bm{k}_{ab}$  = $\bm{K}_{ab}\bm{K}_{bb}^{-1}$, nondimensional unit
$\bm{\Sigma}_{p}$  = process noise covariance, nondimensional unit
$\bm{\Sigma}_{m}$  = measurement noise covariance, nondimensional unit
$\bm{x}_{t}$  = system state at time $t$, nondimensional unit
$\bm{y}_{t}$  = measurement at time $t$, nondimensional unit
$\bm{F}(\cdot)$  = known transition function structure
$f(\cdot)$  = unknown component of the transition function $\bm{F}(\cdot)$
$\bm{g}(\cdot)$  = measurement function
$\bm{f}$  = all function values of $f(\cdot)$, nondimensional unit
$\bm{u}$  = inducing point set, nondimensional unit
$\bm{f}_{\neq u}$  = function values of $f(\cdot)$ excluding the inducing points $\bm{u}$, nondimensional unit
$p(\cdot)$  = exact probability density function
$q(\cdot)$  = approximate probability density function
$\bm{\mu}_{t}$  = mean of the system state $\bm{x}$ at time $t$, nondimensional unit
$\bm{P}_{t}$  = covariance of the system state $\bm{x}$ at time $t$, nondimensional unit
$\bm{m}_{u}$  = mean of the inducing points $\bm{u}$, nondimensional unit
$\bm{V}_{xu}$  = cross-covariance between the system state $\bm{x}$ and the inducing points $\bm{u}$, nondimensional unit
$\bm{S}_{uu}$  = auto-covariance of the inducing points $\bm{u}$, nondimensional unit
$f_{t}$  = function value $f(\bm{\mu}_{t})$, nondimensional unit
$m_{f_{t}}$  = mean of the function value $f_{t}$, nondimensional unit
$\bm{F}_{t}$  = predicted mean of the state transition function, i.e., $\bm{F}_{t}=\bm{F}(\bm{\mu}_{t},m_{f_{t}})$, nondimensional unit
$\bm{A}_{f}$  = transition Jacobian with respect to the function value $f_{t}$, nondimensional unit
$\bm{A}_{x}$  = transition Jacobian with respect to the state $\bm{x}$, nondimensional unit
$\bm{C}_{x}$  = measurement Jacobian with respect to the state $\bm{x}$, nondimensional unit
$\bar{\bm{u}}$  = augmented inducing point set, i.e., $\bar{\bm{u}}=[\bm{u}^{T},f_{t}^{T}]^{T}$, nondimensional unit
$\bm{X}_{t}$  = augmented state at time $t$, i.e., $\bm{X}_{t}=[\bm{x}_{t}^{T},{\bm{u}}^{T}]^{T}$, nondimensional unit
$\bm{\xi}_{t}$  = mean of the augmented state $\bm{X}_{t}$, nondimensional unit
$\bm{\Sigma}_{t}$  = covariance of the augmented state $\bm{X}_{t}$, nondimensional unit
$\bar{\bm{X}}_{t}$  = augmented state with $\bar{\bm{u}}$, i.e., $\bar{\bm{X}}_{t}=[\bm{x}_{t}^{T},\bar{\bm{u}}^{T}]^{T}$, nondimensional unit
$\bar{\bm{\xi}}_{t}$  = mean of the augmented state $\bar{\bm{X}}_{t}$, nondimensional unit
$\bar{\bm{\Sigma}}_{t}$  = covariance of the augmented state $\bar{\bm{X}}_{t}$, nondimensional unit
$\bm{\Phi}$  = transition Jacobian for the augmented state $\bar{\bm{X}}_{t}$, nondimensional unit
$\bm{\Sigma}_{p,\bm{X}}$  = process noise covariance for the augmented state $\bar{\bm{X}}_{t}$, nondimensional unit
$\bm{H}$  = measurement Jacobian for the augmented state $\bar{\bm{X}}_{t}$, nondimensional unit
$M$  = budget for inducing point set size, nondimensional unit
$u_{\mathrm{d}}$  = inducing point to be discarded, nondimensional unit
$\bm{u}_{\mathrm{l}}$  = inducing points left after discarding, nondimensional unit
$i_{\mathrm{d}}$  = index of the inducing point to be discarded
$s_{d}$  = score evaluting the importance of the inducing points, nondimensional unit
$\bm{Q}$  = inversion of the kernel matrix $\bm{K}_{uu}^{-1}$, nondimensional unit
$\bm{\Omega}$  = inversion of the joint covariance $\bm{\Sigma}_{t}^{-1}$, nondimensional unit
$\bm{Q}_{dd}$  = the $i_{\mathrm{d}}$-th diagonal element of the matrix $\bm{Q}$, nondimensional unit
$\bm{Q}_{du}$  = the $i_{\mathrm{d}}$-th row of the matrix $\bm{Q}$, nondimensional unit
$\bm{\Omega}_{dd}$  = the $(d_{x}+i_{\mathrm{d}})$-th diagonal element of the matrix $\bm{\Omega}$, nondimensional unit
$\gamma$  = novelty metric for inducing points, nondimensional unit
$\varepsilon_{\text{tol}}$  = threshold for adding inducing points, nondimensional unit
Superscripts
-  = predicted distribution
*  = optimal distribution
old  = old GP hyperparameters
new  = new GP hyperparameters
@@ 1 Introduction
State-space models (SSMs), also known as hidden Markov models (HMMs), describe the evolution and observation processes of latent states through a state transition model and a measurement model. By summarizing historical information into the latent state, SSMs provide a concise representation of system dynamics and are widely applied in fields such as aerospace engineering [1, 2], target tracking [3, 4, 5], and control systems [6, 7, 8, 9]. Nevertheless, due to the complexity of real-world systems, establishing precise SSMs based on first principles is challenging. To achieve sufficiently accurate system models, extensive research has focused on data-driven SSM modeling [10, 11, 12, 13].
In the past decade, a novel SSM learning method based on the Gaussian process (GP), a nonparametric modeling technique, has been explored [14, 12]. Mathematically, a GP is a probability distribution over functions, providing a flexible and probabilistic framework for modeling unknown functions. GPs not only yield predictions but also quantify uncertainty, making them widely used in machine learning [15] and system identification [16]. Over the past 30 years, many limitations of GPs, including cubic computational complexity, limited representational capacity, and challenges with online learning, have been gradually overcome. For example, various sparse approximation techniques have been developed, such as inducing point methods [17, 18], spectrum approximations [19, 20] and Vecchia approximation [21]. Recent advances in scalable GPs [22, 23] have enabled GPs to handle datasets with millions of points. To improve representational power, deep GPs [24, 25] and ensemble GPs [26] have been developed by stacking multiple layers of GPs and mixing multiple GPs, respectively. Besides, online extensions like sparse online GPs [27], local GPs [28], and streaming sparse GPs [29] have equipped GPs with the ability to process streaming data. Therefore, GPs have gradually become a powerful tool for various applications.
@@
Building on the aforementioned advances, GPs have been naturally developed for the modeling of dynamical systems. In the context of SSM learning, GPs can be employed to model the transition and measurement functions, resulting in what are known as Gaussian process state-space models (GPSSMs) [12]. By employing GPs, GPSSMs offer flexible learning for the dynamical system and provide prediction uncertainty, which is crucial for safety-critical applications [30]. To date, various GPSSM methods have been developed, including sample-based methods [31, 32, 33, 34] and variational inference (VI)-based methods [35, 36, 37, 38, 39]. For example, the early variational GPSSM method proposed by Frigola et al. [32] combines variational Bayes and sequential Monte Carlo (SMC) to learn the GPSSM by iteratively updating the GP and state trajectory. Subsequently, the probabilistic recurrent state space model (PRSSM) [36] adopts the training approach of recurrent neural networks (RNNs) to optimize the variational posterior of the initial state and GP. In view of the limitation of the mean-field assumption in PRSSM, the variationally coupled dynamic trajectories (VCDT) [37] parameterize the coupling between the state and inducing points. Recently, the ensemble Kalman filter-aided variational inference (EnVI) [39] employs the ensemble Kalman filter (EnKF) to capture the coupling between state and GP, eliminating the need for the additional inference network required in previous methods. These methods have achieved significant progress in GPSSM learning, but most studies focus solely on offline scenarios. In engineering practice, many systems experience non-stationary changes (e.g., drone propeller performance varies with battery levels), and there is the possibility of operating in out-of-distribution settings. Therefore, this underscores the urgent need for online learning methods for GPSSMs, which is the focus of this paper.
@@
Facilitating online learning for GPSSMs presents significant challenges. Beyond the usual constraints associated with online learning—such as minimizing memory usage, reducing computational complexity, and preventing catastrophic forgetting—GPSSMs pose additional, unique difficulties. These challenges are detailed as follows. (1) Nonlinearity complicates inference. In GPSSMs, the system state is included in the input to the GP model, introducing inherent nonlinearity. This nonlinearity is intractable, which prevents exact inference through analytic methods. (2) Nonparametric nature of GP. For standard Gaussian process regression (GPR), the training complexity scales cubically with the number of samples [15]. In an online context, as new data arrives continuously, the computational cost of GPs increases rapidly, rendering inference intractable. (3) Coupling between the system state and the system model. Since the state is implicitly measured in general, learning the system model requires inferring the system state. However, state inference depends on the system model itself. As a result, there is a coupling between the state and the model, which increases the complexity of inference. Theoretically, since both are uncertain, they should be treated as latent variables and inferred simultaneously. (4) Hyperparameter online optimization. GPSSMs involve several hyperparameters, such as GP kernel parameters, which impact the learning accuracy. In online settings, hyperparameter optimization presents two main challenges. First, due to memory efficiency considerations, historical measurements are not retained in online scenarios, depriving hyperparameter optimization of explicit information sources. Secondly, since the GP prior depends on kernel hyperparameters, tuning these parameters requires corresponding adjustments to the posterior distribution. Hence, the online GPSSM problem is quite challenging, and as a result, relevant research remains limited.
@@
To overcome the aforementioned shortcomings, a recursive learning algorithm for GPSSMs is proposed in this paper, capable of online adaptation to the operating domain and GP hyperparameters. The main contributions can be summarized as follows: Firstly, to implement a domain-independent recursive learning algorithm for GPSSMs, a novel Bayesian update equation is derived without pre-parameterizing the GP. In the derivation process, only mild linearization approximations are applied to achieve a closed-form solution, offering advantages in learning speed and stability. Through continuously expanding the inducing-point set, this method can accommodate real-time data distributions. Secondly, to address the computational cost issue caused by the increasing number of inducing points, an selection algorithm is developed with both addition and removal operations. To minimize accuracy loss, an optimal discarding equation and selection metric for inducing points are derived based on informative criteria. Thirdly, to alleviate the burden of offline hyperparameter tuning, an online GP hyperparameter optimization method is proposed. In this method, an approximate observation likelihood model is extracted from the obtained posterior distribution. Utilizing this likelihood model as the information source, the GP hyperparameters are optimized, and the posterior distribution can be adjusted as the hyperparameters change. Fourthly, to ensure numerical stability, a stable implementation method based on Cholesky decomposition is developed. Additionally, the extension to multi-output GP is provided, which significantly enhances the practicality and flexibility. Finally, the effectiveness of the proposed method is validated through two synthetic simulations and a real-world dataset. The experimental results demonstrate the superior accuracy, computational efficiency, and adaptability of the proposed method compared to state-of-the-art online GPSSM techniques.
@@ 2.1 Gaussian Processes
where $m_{0}(\cdot)$ denotes the prior mean function and $K_{0}(\cdot,\cdot)$ represents the prior covariance function (also known as kennel function), and $\bm{\theta}$ denotes the kernel hyperparameters. As a common setting, we assume zero-mean GP prior, namely, $m_{0}(\cdot)\equiv 0$. Then, given noise-free function values $\bm{f}_{\mathcal{X}}=\{f(\textbf{x}_{i})\}_{i=1}^{n_{\mathcal{X}}}$ at inputs $\mathcal{X}=\{\mathbf{x}_{i}\}_{i=1}^{n_{\mathcal{X}}}$, the posterior distribution for test function values $\bm{f}_{*}=\{f(\textbf{x}_{i}^{*})\}_{i=1}^{n_{\mathcal{X}^{*}}}$ at inputs $\mathcal{X}^{*}=\{\mathbf{x}_{i}^{*}\}_{i=1}^{n_{\textbf{X}^{*}}}$ is [15]:
@@
Here, we use shorthand notation for the kernel matrix, e.g., $\bm{K}_{**}$ denotes the auto-covariance of $\bm{f}_{*}$, and $\bm{K}_{*\mathcal{X}}$ the cross-covariance between $\bm{f}_{*}$ and $\bm{f}_{\mathcal{X}}$. This notation convention is consistently used throughout the entire text. The above discussion is based on single-output GPs, but indeed GPs can be utilized to learn multi-output functions by employing an associated kernel. However, for simplicity and clarity, the proposed method is elaborated first based on the single-output GP, and the extension to the multi-output case is discussed in Section 5.1. Hence, in the following, the function to learn will be denoted as $f(\cdot)$ without boldface.
2.2 Online Gaussain Processs State Space Models
@@
where the vector $\bm{x}\in\mathbb{R}^{d_{x}}$ and $\bm{y}\in\mathbb{R}^{d_{y}}$ respectively denote the system state and measurement. The vectors $\bm{\omega}_{p}$ and $\bm{\omega}_{m}$ represent the process and measurement noise, with covariances $\bm{\Sigma}_{p}$ and $\bm{\Sigma}_{m}$, respectively. Additionally, the function $\bm{F}:\mathbb{R}^{d_{x}}\to\mathbb{R}^{d_{x}}$ denotes a known function structure for the transition model666 This structure of $F(\cdot)$ can represent various models of interest, such as when only some dimensions of the transition model are unknown, or a discretized continuous-time transition model $F(x_{t},f(x_{t}))=x_{t}+f(x_{t})\Delta t$, where $\Delta t$ is the time step and the unknown component is the time-derivative model $\dot{x}=f(x)$. Through using a varying time step $\Delta t$, the latter can be used for irregular measurement scenarios commonly occurring in biochemical sciences. , which contains an unknown component, represented as the function $f:\mathbb{R}^{d_{x}}\to\mathbb{R}$. To ensure identifiability, the measurement model $\bm{g}:\mathbb{R}^{d_{x}}\to\mathbb{R}^{d_{y}}$ is assumed to be known777 Even if the measurement model is unknown in practice, we can augment the measurement $\bm{y}$ into the state space and transform the problem into one with a known measurement model [12]. .
@@
$f(\cdot)\sim\mathcal{GP}(m_{0}(\cdot),K_{0}(\cdot,\cdot;\bm{\theta}))$ (4b)
$\bm{x}_{t+1}|\bm{x}_{t},\bm{f}\sim\mathcal{N}\left(\bm{x}_{t}\,\big{|}\,\bm{F}(\bm{x}_{t},f(\bm{x}_{t})),\,\bm{\Sigma}_{p}\right)$ (4c)
@@
where $p(\bm{x}_{0})$ denotes a prior Gaussian distribution for the initial state $\bm{x}_{0}$, and we use the notation $\bm{f}$ to denote all function values of the function $f(\cdot)$. For conciseness, we omit the system input from the transition and measurement models, which can be easily incorporated by augmenting it into the input of these models and the function $f(\cdot)$.
In the online setting, measurements $\bm{y}$ arrive sequentially, and we assume that historical measurements are not retained for computational and storage reasons. Given this streaming data setting, the inference task of online GPSSMs is to sequentially approximate the filtering distribution $p(\bm{x}_{t},\bm{f}|\bm{y}_{1:t})$, using the previous result $p(\bm{x}_{t-1},\bm{f}|\bm{y}_{1:t-1})$ and the current measurement $\bm{y}_{t}$. As illustrated in the introduction, several challenges exist in this task, including the nonlinearity and coupling involved, as well as the nonparametric nature of GPs and the hyperparameter optimization problem. To address these difficulties, a novel online GPSSM approach will be developed in the next two sections, with the ability to adapt to both operating domains and hyperparameters.

### v4

@@ Abstract
keywords:
Gaussian process; state-space model; online learning; inducing points; hyperparameter adaptation.
thanks: Corresponding author: Lin Cheng (chenglin5580@buaa.edu.cn).
@@ 1 Introduction
State-space models (SSMs), also known as hidden Markov models (HMMs), describe the evolution and observation processes of latent states through a state transition model and a measurement model. By summarizing historical information into the latent state, SSMs provide a concise representation of system dynamics and are widely applied in fields such as motion planning [1], target tracking [2, 3], and control systems design [4, 5, 6, 7]. Nevertheless, due to the complexity of real-world systems, establishing precise SSMs based on first principles is challenging. To achieve sufficiently accurate system models, extensive research has focused on data-driven SSM modeling [8, 9, 10, 11, 12].
@@
In the past decade, a novel SSM learning method based on the Gaussian process (GP), a nonparametric modeling technique, has been explored [13, 11]. Mathematically, a GP is a probability distribution over the function space, providing a flexible and probabilistic framework for modeling unknown functions. GPs not only yield predictions but also quantify uncertainty, making them widely used in machine learning [14] and system identification [15, 16]. In the context of SSM learning, GPs can be employed to model the transition and measurement functions, resulting in what are known as Gaussian process state-space models (GPSSMs) [11]. By employing GPs, GPSSMs offer flexible learning for the dynamical system and provide prediction uncertainty, which is crucial for safety-critical applications [17]. To date, various GPSSM methods have been developed, including sample-based methods [18, 19, 20, 21] and variational inference (VI)-based methods [22, 23, 24, 25, 26]. For example, the early variational GPSSM method proposed by Frigola et al. [19] combines variational Bayes and sequential Monte Carlo (SMC) to learn the GPSSM by iteratively updating the GP and the state trajectory. Subsequently, the probabilistic recurrent state space model (PRSSM) [23] adopts the training approach of recurrent neural networks (RNNs) to optimize the variational posterior of the initial state and GP. In view of the limitation of the mean-field assumption in PRSSM, the variationally coupled dynamic trajectories (VCDT) [24] parameterize the coupling between the state and inducing points. Recently, the ensemble Kalman filter-aided variational inference (EnVI) [26] employs the ensemble Kalman filter (EnKF) to capture the coupling between state and GP, eliminating the need for the additional inference network required in previous methods. These methods have achieved significant progress in GPSSM learning, but most studies focus solely on offline scenarios. In engineering practice, many systems experience non-stationary changes or varies cross tasks, and there is the possibility of operating in out-of-distribution settings. Therefore, this underscores the urgent need for online learning methods for GPSSMs, which is the focus of this paper.
@@
Facilitating online learning for GPSSMs presents significant challenges. In addition to the typical constraints of online learning, such as minimizing memory usage, reducing computational complexity, and preventing catastrophic forgetting, GPSSMs introduce further unique difficulties. These challenges are described as follows. (1) Nonlinearity complicates inference. In GPSSMs, the system state is included in the input to the GP model, introducing inherent nonlinearity. This nonlinearity is intractable, which prevents exact inference through analytic methods. (2) Nonparametric nature of GP. For standard Gaussian process regression (GPR), the training complexity scales cubically with the number of samples [14]. In an online context, as new data arrives continuously, the computational cost of GPs increases rapidly, rendering inference intractable. (3) Coupling between the system state and the system model. Since the state is implicitly measured in general, learning the system model requires estimating the system state. However, state estimation depends on the system model itself. As a result, there is a coupling between the state and the model, which increases the complexity of inference. Theoretically, since both are uncertain, they should be treated as latent variables and inferred simultaneously. (4) Hyperparameter online optimization. GPSSMs involve several hyperparameters, such as GP kernel parameters, which impact the learning accuracy. In online settings, hyperparameter optimization presents two main challenges. First, due to memory efficiency considerations, historical measurements are not retained in online scenarios, depriving hyperparameter optimization of explicit information sources. Secondly, since the GP prior depends on kernel hyperparameters, tuning these parameters requires corresponding adjustments to the posterior distribution. In summary, the online GPSSM problem is quite challenging, and as a result, relevant research remains limited.
@@
To address the aforementioned shortcomings, this paper proposes a recursive learning algorithm for GPSSMs that is capable of online adaptation to the operating domain and GP hyperparameters. The main contributions can be summarized as follows: Firstly, to implement a domain-independent recursive learning algorithm for GPSSMs, a novel Bayesian update equation is derived without pre-parameterizing the GP. In the derivation process, only mild linearization approximations are applied to achieve a closed-form solution, offering advantages in learning speed and stability. Through continuously expanding the inducing-point set, this method can accommodate real-time data distributions. Secondly, to address the computational cost issue caused by the increasing number of inducing points, a selection algorithm is developed with both addition and removal operations. To minimize accuracy loss, an optimal discarding equation and selection metric for inducing points are derived based on informative criteria. Thirdly, to alleviate the burden of offline hyperparameter tuning, an online GP hyperparameter optimization method is proposed. In this method, an approximate observation likelihood model is extracted from the obtained posterior distribution. Utilizing this likelihood model as the information source, the GP hyperparameters are optimized, and the posterior distribution can be adjusted as the hyperparameters change. Finally, the effectiveness of the proposed method is validated through two synthetic simulations and a real-world dataset. The experimental results demonstrate the superior accuracy, computational efficiency, and adaptability of the proposed method compared to state-of-the-art online GPSSM techniques.
@@ 2.1 Gaussian Processes
where $m_{0}(\cdot)$ denotes the prior mean function and $K_{0}(\cdot,\cdot)$ represents the prior covariance function (also known as kernel function), and $\bm{\theta}$ denotes the kernel hyperparameters. As a common setting, we assume zero-mean GP prior, namely, $m_{0}(\cdot)\equiv 0$. Then, given noise-free function values $\bm{f}_{\mathcal{X}}=\{f(\textbf{x}_{i})\}_{i=1}^{n_{\mathcal{X}}}$ at inputs $\mathcal{X}=\{\mathbf{x}_{i}\}_{i=1}^{n_{\mathcal{X}}}$, the posterior distribution for test function values $\bm{f}_{*}=\{f(\textbf{x}_{i}^{*})\}_{i=1}^{n_{\mathcal{X}^{*}}}$ at inputs $\mathcal{X}^{*}=\{\mathbf{x}_{i}^{*}\}_{i=1}^{n_{\mathcal{X}^{*}}}$ is [14]:
@@
Here, we use shorthand notation for the kernel matrix, e.g., $\bm{K}_{**}$ denotes the auto-covariance of $\bm{f}_{*}$, and $\bm{K}_{*\mathcal{X}}$ the cross-covariance between $\bm{f}_{*}$ and $\bm{f}_{\mathcal{X}}$. This notation convention is consistently used throughout the entire text. The above discussion is based on single-output GPs. However, GPs can also be applied to learn multi-output functions by employing a matrix-valued kernel. However, for simplicity and clarity, the proposed method is elaborated first based on the single-output GP, and the extension to the multi-output case is discussed in Section 5.1. Hence, in the following, the function to learn will be denoted as $f(\cdot)$ without boldface.
2.2 Online Gaussian Process State-Space Models
@@
where the vector $\bm{x}\in\mathbb{R}^{d_{x}}$ and $\bm{y}\in\mathbb{R}^{d_{y}}$ respectively denote the system state and measurement. The vectors $\bm{\omega}_{p}$ and $\bm{\omega}_{m}$ represent the process and measurement noise, with covariances $\bm{\Sigma}_{p}$ and $\bm{\Sigma}_{m}$, respectively. Additionally, the function $\bm{F}:\mathbb{R}^{d_{x}}\times\mathbb{R}\to\mathbb{R}^{d_{x}}$ denotes a known function structure for the transition model [^1] This structure of $\bm{F}(\cdot)$ can represent various models of interest, such as when only some dimensions of the transition model are unknown, or a discretized continuous-time transition model $\bm{F}(\bm{x}_{t},\bm{f}(\bm{x}_{t}))=\bm{x}_{t}+\bm{f}(\bm{x}_{t})\Delta t$, where $\Delta t$ is the time step and the unknown component is the time-derivative model $\dot{\bm{x}}=\bm{f}(\bm{x})$. Through using a varying time step $\Delta t$, the latter can be used for irregular measurement scenarios commonly occurring in biochemical sciences. , which contains an unknown component, represented as the function $f:\mathbb{R}^{d_{x}}\to\mathbb{R}$. To ensure identifiability [26], the measurement model $\bm{g}:\mathbb{R}^{d_{x}}\to\mathbb{R}^{d_{y}}$ is assumed to be known [^2] Even if the measurement model is unknown in practice, we can augment the measurement $\bm{y}$ into the state space and transform the problem into one with a known measurement model [11]. .
@@
$f(\cdot)\sim\mathcal{GP}(\bm{0},K_{0}(\cdot,\cdot;\bm{\theta}))$ (4b)
$\bm{x}_{t+1}|\bm{x}_{t},\bm{f}\sim\mathcal{N}\left(\bm{x}_{t+1}\,\big|\,\bm{F}(\bm{x}_{t},f(\bm{x}_{t})),\,\bm{\Sigma}_{p}\right)$ (4c)
@@
where $p(\bm{x}_{0})=\mathcal{N}(\bm{x}_{0}|\bm{\mu}_{0},\bm{P}_{0})$ denotes a prior Gaussian distribution for the initial state $\bm{x}_{0}$, and we use the notation $\bm{f}=\{f(\bm{x}):\bm{x}\in\mathbb{R}^{d_{x}}\}$ to denote all function values of the function $f(\cdot)$. For conciseness, we omit the control input $\bm{c}\in\mathbb{R}^{d_{c}}$ from the transition and measurement models, which can be easily incorporated by augmenting it into the input of these models and the function $f(\cdot)$.
In the online setting, measurements $\bm{y}$ arrive sequentially, and we assume that historical measurements are not retained for computational and storage reasons. Given this streaming data setting, the inference task of online GPSSMs is to sequentially approximate the filtering distribution $p(\bm{x}_{t},\bm{f}|\bm{y}_{1:t})$, using the previous result $p(\bm{x}_{t-1},\bm{f}|\bm{y}_{1:t-1})$ and the current measurement $\bm{y}_{t}$. As illustrated in the introduction, several challenges exist in this task, including nonlinearity, coupling between state and GP, as well as the nonparametric nature of GPs and the hyperparameter optimization problem. To address these difficulties, a novel online GPSSM approach will be developed in the next two sections, with the ability to adapt to both operating domains and hyperparameters.